2a-3b=-8
4a+3b=-34
I am supposed to be using either substitution, elimination, or inverse matrices to solve.
Thanks in advance for your time and help!

Answer by stanbon(75887)

2a-3b=-8
4a+3b=-34
-----------
Add to eliminate the b-terms:
---
6a = -42
---
a = -7
---
Solve for "b":
2a-3b=-8
2(-7)-3b = -8
-14-3b = -8
Add 14 to both sides to get:
-3b = 6
b = -2
==============
cheers,
Stan H.